# Optical design and sensitivity optimization of Cryogenic sub-Hz cROss torsion bar detector with quantum NOn-demolition Speed meter (CHRONOS)

### v1

@@ VII Discussion
Realizing such flexibility requires auxiliary optical systems in addition to the main interferometer. In particular, a dedicated alignment system using a green laser and wavefront sensors (WFS) is necessary. Given the relatively low transmissivity of the PRC input mirror, such systems are essential for maintaining mode matching and cavity stability. These subsystems are also crucial for ensuring robust operation and reliable calibration in the presence of environmental disturbances.
It must also be noted that our discussion is restricted to the 2.5 m laboratory-scale facility. While this configuration allows detailed investigation of quantum-noise shaping and control techniques, scaling the arm length inevitably alters the balance of noise contributions. Among these, coating thermal noise remains a decisive factor. Even at the 2.5-m scale, the choice of coating technology—mechanical loss, thickness, and deposition method— directly determines the mid-band sensitivity floor Harry et al. (2002); Levin (1998). Future advances in low-loss coating materials will therefore be indispensable for realizing the scientific goals of CHRONOS and extending the design to larger-scale detectors.
@@
In conclusion, the optical design and sensitivity modeling presented here establish the foundation of CHRONOS as a versatile sub-hertz interferometer. Although its scale is limited, the 2.5-m system highlights key trade-offs in cavity detuning, detection quadrature, and coating technology, providing valuable guidance for both laboratory-scale research and the path toward future long-baseline detectors.
@@ VIII Conclusion
The optimized configuration predicts a quantum-noise-limited strain sensitivity of
$h\simeq 3\times 10^{-18}\penalty 10000\ \mathrm{Hz^{-1/2}}\quad\text{at }1\penalty 10000\ \mathrm{Hz},$ (25)
with a ring-cavity finesse of $\mathcal{F}\simeq 3.1\times 10^{4}$ and a round-trip Gouy phase of $\psi_{\mathrm{rt}}\approx 153^{\circ}$. The power-recycling cavity (PRC) detuning, optimized at $\phi_{p}=-85^{\circ}$, was identified as the dominant factor shaping the low-frequency quantum noise, while the signal-recycling cavity (SRC) detuning ($\phi_{s}=0^{\circ}$) introduces a uniform quadrature rotation around $90^{\circ}$ without improving sensitivity. The optimal homodyne detection angle was determined to be $\zeta_{\mathrm{opt}}\simeq 46^{\circ}$, where the trade-off between radiation-pressure suppression and shot-noise minimization yields the best overall sensitivity near 1 Hz.
For the end test masses (ETMs), a target reflectivity of $R_{\mathrm{ETM}}=99.9999\%$ is assumed. However, optical absorption in the coatings remains the dominant limitation to the overall sensitivity, particularly under cryogenic operation around 10 K. Further development of low-absorption dielectric materials such as SiN and SiON will therefore be indispensable for extending CHRONOS toward next-generation cryogenic detectors.
@@
The 2.5 m CHRONOS interferometer thus serves as both a versatile testbed for precision quantum-noise studies and a practical platform for environmental-noise investigations. The system exhibits high optical stability ($\psi_{\mathrm{rt}}\approx 153^{\circ}$), large finesse ($\mathcal{F}\approx 3\times 10^{4}$), and sub-hertz quantum-limited sensitivity of $3\times 10^{-18}\penalty 10000\ \mathrm{Hz^{-1/2}}$ at 1 Hz. These results establish a solid foundation for scaling CHRONOS toward longer-baseline facilities, bridging laboratory-scale quantum optics with the development of next-generation sub-hertz gravitational-wave observatories.
@@ Acknowledgements.
We also thank Kin-Wang Ng and Masashi Hazumi for their academic advice during the preparation of this manuscript. We thank Masaya Hasegawa, Takahiro Kanayama, Satoki Matsushita, Hirokazu Murakami, Mai Inoue, Ryuji Shibuya and Chia-Ming Kuo for support to establish CHRONOS team. Y.I. and H.W. acknowledges support from NSTC, CHiP and Academia Sinica in Taiwan under Grant No.114-2112-M-008-006-, and No.AS-TP-112-M01.
@@ Appendix A Complete Optical Element Parameters
Table 6 lists the full set of optical elements used in the 2.5 m CHRONOS interferometer, including their coordinates, incidence angles, curvature radii, and reflectivities. For brevity, power transmissivity and absorption values are omitted; they can be obtained from the relation T=1RAT=1-R-A, with $A\approx 0$.
Table 6: Full list of optical elements in the CHRONOS 2.5 m. The corresponding optical components are shown in Fig. 2. Incident angle in degrees, curvature radius in meters, and power reflectivity $R$. Based on the length of Ring cavity, PRC, and SRC, the optimized radius of curvature is listed.
@@ Appendix B Noise Model
This appendix summarizes the principal technical noise models included in the sensitivity analysis. All contributions are normalized to the detector output using the frequency-dependent sensitivity transfer function $H_{\mathrm{sens}}(f)$ and the arm length $L_{\mathrm{arm}}$.
@@ B.1 Coating Brownian noise
Brownian motion arising from the dielectric coating of the ETMs is one of the dominant fundamental noise sources in the mid-frequency band of CHRONOS. This arises from microscopic mechanical dissipation in the high-reflectivity multilayer coating, which drives thermally induced surface displacement through the fluctuation-dissipation theorem Levin (1998); Harry et al. (2002). In the CHRONOS design, the interferometer response is ultimately limited by this process at frequencies where seismic and suspension noise are already suppressed and radiation-pressure noise is no longer dominant.
@@
$S_{\mathrm{coat}}(f)$ $=\frac{1}{L_{\mathrm{arm}}}\sqrt{\frac{2k_{\mathrm{B}}T}{\pi^{2}f}}\sqrt{\frac{1-\sigma_{s}^{2}}{\omega_{\mathrm{ETM}}Y_{s}}}\sqrt{\phi_{c}^{\mathrm{eff}}}.$ (26)
Here:
  • $L_{\mathrm{arm}}$ is the effective arm length of the torsion bar. The prefactor $1/L_{\mathrm{arm}}$ converts the strain noise of the ETM surface into an equivalent strain noise.
  • $k_{\mathrm{B}}$ is Boltzmann’s constant, and $T$ is the temperature of the coating (cryogenic in the CHRONOS configuration). The scaling $\propto\sqrt{T}$ explicitly shows why cryogenic operation is essential for suppressing coating Brownian noise.
  • $\sigma_{s}$ and $Y_{s}$ are the Poisson ratio and Young’s modulus of the ETM substrate, respectively. These elastic parameters determine how efficiently internal mechanical loss in the coating is converted into surface motion.
  • $\omega_{\mathrm{ETM}}$ is a geometric factor set by the spatial profile of the sensing beam on the mirror surface. For a Gaussian interrogation beam, $\omega_{\mathrm{ETM}}$ is approximately proportional to the beam radius $w$ on the ETM. A larger beam radius averages over a wider area and thereby reduces the coupling to local coating fluctuations.
  • $\phi_{c}^{\mathrm{eff}}$ is the effective mechanical loss angle of the coating stack. This term incorporates the individual loss angles, thicknesses, and elastic moduli of the high-index and low-index layers in the dielectric mirror coating. Multilayer designs with low-loss (e.g., SiN/SiON) materials reduce $\phi_{c}^{\mathrm{eff}}$ and thus directly improve sensitivity.
The $f^{-1/2}$ scaling of $S_{\mathrm{coat}}(f)$ reflects the thermally driven random force spectrum acting on a mechanically lossy elastic system. In CHRONOS, this contribution sets the practical noise floor in the mid-frequency range (roughly $\sim 1$$10\penalty 10000\ \mathrm{Hz}$), where radiation-pressure noise has begun to decouple due to the speed-meter topology, but seismic and suspension noise are already strongly attenuated. Consequently, coating Brownian noise defines one of the key material-driven limits to the interferometer performance and directly motivates the use of cryogenic, low-loss coating technologies (Sec. A and Sec. VI.1).
B.2 Bar Thermal Noise (Torsion Bar)
The torsion-bar thermal noise model combines two principal contributions: (1) the broadband structural damping arising from internal material loss of the suspension and bar body, and (2) the narrow-band angular thermal noise peaking near the torsional resonance frequency Saulson (1990); Yamamoto et al. (2006). These two effects together describe both the wideband background and the resonant thermal motion of the torsion bar.
@@ B.2 Bar Thermal Noise (Torsion Bar)
$S_{\mathrm{bar,bg}}(f)$ $=\sqrt{\frac{2k_{\mathrm{B}}T}{\pi^{3/2}f}}\sqrt{\frac{1-\sigma_{s}^{2}}{\omega_{\mathrm{ETM}}Y_{s}}}\sqrt{\phi_{s}^{\mathrm{eff}}},$ (27)
$S_{\mathrm{bar,rot}}(f)$ $=\sqrt{\frac{4k_{\mathrm{B}}T}{I}}\sqrt{\frac{\eta^{2}\omega_{\mathrm{tb}}\phi_{0}}{(\omega_{\mathrm{tb}}^{2}-\omega^{2})^{2}+\omega\omega_{\mathrm{tb}}\phi_{0}}},$ (28)
$S_{\mathrm{bar}}(f)$ $=\frac{1}{L_{\mathrm{arm}}}\sqrt{S_{\mathrm{bar,bg}}^{2}+S_{\mathrm{bar,rot}}^{2}}.$ (29)
Here $\omega=2\pi f$, $\omega_{\mathrm{tb}}$ is the torsional resonance angular frequency, $\phi_{s}^{\mathrm{eff}}$ is the effective material loss angle, $\phi_{0}$ represents the loss parameter associated with the resonance peak, $\eta$ denotes the coupling efficiency between the bar rotation and interferometer readout, and $I$ is the moment of inertia of the torsion bar.
The first term, $S_{\mathrm{bar,bg}}$, represents the broadband contribution arising from structural damping and internal friction of the suspension fibers or bar material. It exhibits a characteristic $f^{-1/2}$ dependence typical of thermally driven losses. The second term, $S_{\mathrm{bar,rot}}$, describes the resonant enhancement near $\omega\approx\omega_{\mathrm{tb}}$, reflecting the thermal excitation of the torsional eigenmode. At frequencies well below resonance, $S_{\mathrm{bar,rot}}$ scales approximately as $f^{-1}$, while above resonance it decreases as $f^{-3}$, consistent with the mechanical response of a torsional oscillator.
The total contribution $S_{\mathrm{bar}}$ is converted to strain noise by dividing by the torsion bar arm length $L_{\mathrm{arm}}$. In CHRONOS, this term dominates below $\sim 0.5\penalty 10000\ \mathrm{Hz}$, where seismic noise and radiation-pressure noise are both suppressed. Minimizing the effective loss angles $\phi_{s}^{\mathrm{eff}}$ and $\phi_{0}$, and maximizing the mechanical quality factor $Q=1/\phi_{0}$, are therefore essential design goals for achieving the sub-hertz sensitivity required for CHRONOS.
@@ B.3 Seismic Noise
Noise due to ground displacement couples into the detector output through the pre-isolation system and the suspension transfer functions Saulson (1984); Harms (2019). This coupling is especially critical in the sub-hertz band, where seismic motion dominates all other technical noise sources. In this study, the ground displacement spectrum measured at the ASGRAF ASG site is adopted as the baseline input, combined with the modeled attenuation provided by the pre-isolation stage and the suspension system.
At low frequencies ($f<1\penalty 10000\ \mathrm{Hz}$), the system is assumed to employ a pre-isolation configuration analogous to that of LIGO, incorporating both passive and active isolation stages, followed by a multi-stage pendulum suspension. The overall transfer function thus includes the full yaw response of the suspension. For the rotational degree of freedom, the effective seismic coupling is empirically found to be roughly $1/100$ of the translational coupling amplitude, consistent with the TOBA analysis Ando et al. (2010).
@@
$S_{\mathrm{seis}}(f)=\frac{1}{\omega}\frac{|H_{\mathrm{yaw}}(f)|}{G_{\mu g}G_{\mathrm{pre}}}S_{x}^{\mathrm{ground}}(f),$ (30)
where:
  • $S_{x}^{\mathrm{ground}}(f)$ is the measured ground-displacement spectrum at the ASGRAF site, typically characterized by $10^{-7}$$10^{-9}\penalty 10000\ \mathrm{m/\sqrt{Hz}}$ in the 0.1–10 Hz range.
  • $H_{\mathrm{yaw}}(f)$ is the suspension yaw transfer function, describing the angular response of the suspended bar to ground motion.
  • $G_{\mu g}$ is the conversion gain between mechanical torque and equivalent gravitational acceleration, incorporating the bar geometry and lever arm.
  • $G_{\mathrm{pre}}$ represents the frequency-dependent attenuation of the pre-isolation system, typically providing suppression proportional to $f^{-4}$ to $f^{-6}$ above the microseismic peak.
@@
For the full CHRONOS facility, a dedicated vibration-isolation platform and cryogenic suspension are planned to achieve enhanced attenuation in the sub-hertz region. Such systems will combine multiple layers of active inertial sensing, low-frequency tilt decoupling, and cryogenic damping control to suppress seismic noise below $10^{-18}\penalty 10000\ \mathrm{m/\sqrt{Hz}}$ at frequencies approaching $0.1\penalty 10000\ \mathrm{Hz}$. These developments are essential for extending the CHRONOS sensitivity into the deep sub-hertz band, where Newtonian noise and coating thermal noise become the dominant limiting factors.

### v2

@@ VII Discussion
Realizing such flexibility requires auxiliary optical systems in addition to the main interferometer. In particular, a dedicated alignment system using a green laser and wavefront sensors is necessary. Given the relatively low transmissivity of the PRC input mirror, such systems are essential for maintaining mode matching and cavity stability. These subsystems are also crucial for ensuring robust operation and reliable calibration in the presence of environmental disturbances.
It must also be noted that the discussion presented here is restricted to the 2.5 m laboratory-scale facility. While this configuration allows detailed investigation of quantum-noise shaping and control techniques, scaling the arm length inevitably alters the balance of noise contributions. Among these, coating thermal noise remains a decisive factor. Even at the 2.5 m scale, the choice of coating technology—mechanical loss, thickness, and deposition method— directly determines the mid-band sensitivity floor Harry2002 ; Levin1998 . Future advances in low-loss coating materials will therefore be indispensable for realizing the scientific goals of CHRONOS and extending the design to larger-scale detectors.
@@
In conclusion, the optical design and sensitivity modeling presented here establish the foundation of CHRONOS as a versatile sub-hertz interferometer. Although its scale is limited, the 2.5 m system highlights key trade-offs in cavity detuning, detection quadrature, and coating technology, providing valuable guidance for both laboratory-scale research and the path toward future long-baseline detectors.
@@ VIII Conclusion
With the optimized optical parameters identified in this study, we find that the interferometer can operate at the quantum-noise limit, reaching a strain sensitivity of
$h\simeq 3\times 10^{-18}\penalty 10000\ \mathrm{Hz^{-1/2}}\quad\text{at }1\penalty 10000\ \mathrm{Hz}.$ (30)
This performance is achieved with a ring-cavity finesse of $\mathcal{F}\simeq 3.1\times 10^{4}$ and a round-trip Gouy phase of $\psi\approx 153^{\circ}$, which together define a stable optical eigenmode suitable for low-frequency operation.
Our analysis shows that the detuning of the power-recycling cavity plays a central role in determining the low-frequency quantum-noise behavior. In particular, an optimized PRC detuning of $\phi_{p}=-85^{\circ}$ significantly suppresses radiation-pressure noise and sets the overall noise floor. In contrast, the signal-recycling cavity detuning ($\phi_{s}=0^{\circ}$) primarily results in a nearly uniform quadrature rotation around $90^{\circ}$ and does not lead to a net improvement in sensitivity.
Finally, we identify an optimal homodyne detection angle of $\zeta\simeq 46^{\circ}$, at which the competing requirements of radiation-pressure-noise reduction and shot-noise suppression are best balanced. This choice maximizes the interferometer sensitivity in the vicinity of 1 Hz and highlights the importance of jointly optimizing cavity detuning and readout for low-frequency, quantum-noise-limited interferometers.
For the ETMs, a target reflectivity of $R_{\mathrm{ETM}}=99.9999\%$ is assumed. However, optical absorption in the coatings remains the dominant limitation to the overall sensitivity, particularly under cryogenic operation around 10 K. Further development of low-absorption dielectric materials such as SiN and SiON will therefore be indispensable for extending CHRONOS toward next-generation cryogenic detectors.
@@
The 2.5 m CHRONOS interferometer thus serves as both a versatile testbed for precision quantum-noise studies and a practical platform for environmental-noise investigations. The system exhibits high optical stability ($\psi\approx 153^{\circ}$), large finesse ($\mathcal{F}\approx 3\times 10^{4}$), and sub-hertz quantum-limited sensitivity of $3\times 10^{-18}\penalty 10000\ \mathrm{Hz^{-1/2}}$ at 1 Hz. These results establish a solid foundation for scaling CHRONOS toward longer-baseline facilities, bridging laboratory-scale quantum optics with the development of next-generation sub-hertz gravitational-wave observatories.
@@ Acknowledgements.
We also thank Kin-Wang Ng and Masashi Hazumi for their academic advice during the preparation of this manuscript. We thank Masaya Hasegawa, Takahiro Kanayama, Satoki Matsushita, Hirokazu Murakami, Mai Inoue, Ryuji Shibuya, and Chia-Ming Kuo for their support in establishing the CHRONOS team. Y.I. is supported by NSTC, CHiP, and Academia Sinica in Taiwan under Grant Nos. 114-2112-M-008-006- and AS-TP-112-M01.
Conflict of interest statement
The authors declare no conflicts of interest.
Ethics statement
This research does not involve human participants, human data, or animals.
@@ Appendix A Complete Optical Element Parameters
Table 6 lists the full set of optical elements used in the 2.5 m CHRONOS interferometer, including their coordinates, incidence angles, radii of curvature, and reflectivities. For brevity, power transmissivity and absorption values are omitted; they can be obtained from the relation T=1RAT=1-R-A, with $A\approx 0$.
Table 6: Full list of optical elements in the CHRONOS 2.5 m interferometer. The corresponding optical components are shown in Fig. 2. Incident angles are given in degrees, curvature radii in meters, and power reflectivity as $R$. Based on the lengths of the ring cavity, PRC, and SRC, the optimized radii of curvature are listed.
@@ Appendix B Noise Model
This appendix summarizes the principal technical noise models included in the sensitivity analysis. All contributions are normalized to the detector output using the frequency-dependent sensing transfer function $H_{\mathrm{sens}}(f)$ and the arm length $L_{\mathrm{arm}}$. The assumed parameters are listed in Table 7.
@@ B.1 Coating Brownian noise
Brownian motion arising from the dielectric coating of the ETMs is one of the dominant fundamental noise sources in the mid-frequency band of CHRONOS. This arises from microscopic mechanical dissipation in the high-reflectivity multilayer coating, which drives thermally induced surface displacement through the fluctuation-dissipation theorem Levin1998 ; Harry2002 . In the CHRONOS design, the interferometer response is ultimately limited by this process at frequencies where seismic and suspension noise are already suppressed and radiation-pressure noise is no longer dominant.
@@
$S_{\mathrm{coat}}(\Omega)$ $=\frac{\eta_{g}}{L_{\mathrm{bar}}}\sqrt{\frac{4k_{\mathrm{B}}T}{\pi\Omega}}\sqrt{\frac{1-\sigma_{s}^{2}}{\omega_{\mathrm{ETM}}Y_{s}}}\sqrt{\phi_{c}^{\mathrm{eff}}}.$ (31)
Here, the effective arm length of the torsion bar is defined as $L_{\mathrm{arm}}$. The prefactor $\eta_{g}/L_{\mathrm{arm}}$ converts the displacement noise of the ETM surface into an equivalent strain noise. Boltzmann’s constant is denoted by $k_{\mathrm{B}}$, and the coating temperature is denoted by $T$. The scaling $\propto\sqrt{T}$ explicitly shows why cryogenic operation is essential for suppressing coating Brownian noise. The Poisson ratio and Young’s modulus of the ETM substrate are denoted by $\sigma_{s}$ and $Y_{s}$, respectively. These elastic parameters determine how efficiently internal mechanical loss in the coating is converted into surface motion. The beam width set by the spatial profile of the sensing beam on the mirror surface is denoted by $\omega_{\mathrm{ETM}}$. The effective mechanical loss angle of the coating stack is denoted by $\phi_{c}^{\mathrm{eff}}$. This term incorporates the individual loss angles, thicknesses, and elastic moduli of the high-index and low-index layers in the dielectric mirror coating. Multilayer designs with low-loss (e.g., SiN/SiON) materials reduce $\phi_{c}^{\mathrm{eff}}$ and thus directly improve sensitivity.
The $f^{-1/2}$ scaling of $S_{\mathrm{coat}}(f)$ reflects the thermally driven random-force spectrum acting on a mechanically lossy elastic system. In CHRONOS, this contribution sets the practical noise floor in the mid-frequency range (roughly $\sim 1$$10\penalty 10000\ \mathrm{Hz}$), where radiation-pressure noise has begun to decouple due to the speed-meter topology, while seismic and suspension noise are already strongly attenuated. Consequently, coating Brownian noise defines one of the key material-driven limits to interferometer performance and directly motivates the use of cryogenic, low-loss coating technologies.
B.2 Bar Thermal Noise
The torsion-bar thermal noise model combines two principal contributions: (1) the broadband structural damping arising from internal material loss of the suspension and bar body, and (2) the narrow-band angular thermal noise peaking near the torsional resonance frequency Saulson1990 ; Yamamoto2006 . These two effects together describe both the wideband background and the resonant thermal motion of the torsion bar.
@@ B.2 Bar Thermal Noise
$S_{\mathrm{bar,bg}}(\Omega)$ $=\sqrt{\frac{4k_{\mathrm{B}}T}{\pi^{1/2}\Omega}}\sqrt{\frac{1-\sigma_{s}^{2}}{\omega_{\mathrm{ETM}}Y_{s}}}\sqrt{\phi_{s}^{\mathrm{eff}}},$ (32)
$S_{\mathrm{bar,rot}}(\Omega)$ $=\sqrt{\frac{4k_{\mathrm{B}}T}{I}}\sqrt{\frac{\eta^{2}\omega_{\mathrm{tb}}\phi_{0}}{(\omega_{\mathrm{tb}}^{2}-\Omega^{2})^{2}+\Omega\omega_{\mathrm{tb}}\phi_{0}}},$ (33)
$S_{\mathrm{bar}}(\Omega)$ $=\frac{\eta_{g}}{L_{\mathrm{bar}}}\sqrt{S_{\mathrm{bar,bg}}^{2}(\Omega)+S_{\mathrm{bar,rot}}^{2}(\Omega)},$ (34)
where the torsional resonance angular frequency is defined as $\omega_{\mathrm{tb}}$, $\phi_{s}^{\mathrm{eff}}$ is the effective material loss angle, $\phi_{0}$ represents the loss parameter associated with the resonance peak, $\eta_{g}$ denotes the coupling efficiency between the bar rotation and interferometer readout, and $I$ is the moment of inertia of the torsion bar.
The first term, $S_{\mathrm{bar,bg}}(\Omega)$, represents the broadband contribution arising from structural damping and internal friction of the suspension fibers or bar material. It exhibits a characteristic $\Omega^{-1/2}$ dependence typical of thermally driven losses. The second term, $S_{\mathrm{bar,rot}}$, describes the resonant enhancement near $\omega\approx\omega_{\mathrm{tb}}$, reflecting the thermal excitation of the torsional eigenmode. At frequencies well below resonance, $S_{\mathrm{bar,rot}}$ scales approximately as $\Omega^{-1}$, while above resonance it decreases as $\Omega^{-3}$, consistent with the mechanical response of a torsional oscillator.
The total contribution $S_{\mathrm{bar}}(\Omega)$ is converted to strain noise by dividing by the torsion-bar arm length $L_{\mathrm{arm}}$. In CHRONOS, this term dominates below $\sim 0.5\penalty 10000\ \mathrm{Hz}$, where seismic noise and radiation-pressure noise are both suppressed. Minimizing the effective loss angles $\phi_{s}^{\mathrm{eff}}$ and $\phi_{0}$, and maximizing the mechanical quality factor $Q=1/\phi_{0}$, are therefore essential design goals for achieving the sub-hertz sensitivity required for CHRONOS.
@@ B.3 Seismic Noise
Noise due to ground displacement couples into the detector output through the pre-isolation system and the suspension transfer functions Saulson1984 ; Harms2019 . This coupling is especially critical in the sub-hertz band, where seismic motion dominates all other technical noise sources. In this study, the ground-displacement spectrum measured at the ASGRAF ASGRAF site is adopted as the baseline input, combined with the modeled attenuation provided by the pre-isolation stage and the suspension system.
At low frequencies ($f<1\penalty 10000\ \mathrm{Hz}$), the system is assumed to employ a pre-isolation configuration analogous to that of LIGO, incorporating both passive and active isolation stages, followed by a multi-stage pendulum suspension. The overall transfer function thus includes the full yaw response of the suspension. For the rotational degree of freedom, the effective seismic coupling is empirically found to be roughly $1/100$ of the translational coupling amplitude, consistent with the TOBA analysis Ando2010 .
@@
$S_{\mathrm{seis}}(\Omega)=\frac{|H_{\mathrm{yaw}}(\Omega)|}{G_{\mu g}\,G_{\mathrm{pre}}}S_{x}^{\mathrm{ground}}(\Omega),$ (35)
@@
where $S_{x}^{\mathrm{ground}}(\Omega)$ denotes the ground-displacement spectrum. In this work, $S_{x}^{\mathrm{ground}}(\Omega)$ is calculated assuming measured underground seismic data at the KAGRA site, as reported by Michimura et al. Michimura_2017 . The transfer function $H_{\mathrm{yaw}}(\Omega)$ represents the yaw response of the suspended torsion bar to ground motion. The factor $G_{\mu g}$ is the conversion gain between mechanical torque and the equivalent gravitational acceleration, which incorporates the bar geometry and the effective lever arm. The term $G_{\mathrm{pre}}$ describes the frequency-dependent attenuation of the pre-isolation system, typically providing suppression proportional to $\Omega^{-4}$ to $\Omega^{-6}$ above the microseismic peak. For the full CHRONOS facility, a dedicated vibration-isolation platform and cryogenic suspension are planned to achieve enhanced attenuation in the sub-hertz region. Such systems will combine multiple layers of active inertial sensing, low-frequency tilt decoupling, and cryogenic damping control to suppress seismic noise below $10^{-18}\penalty 10000\ \mathrm{m/\sqrt{Hz}}$ at frequencies approaching $0.1\penalty 10000\ \mathrm{Hz}$. These developments are essential for extending the CHRONOS sensitivity into the deep sub-hertz band, where Newtonian noise and coating thermal noise become the dominant limiting factors.
B.4 Rayleigh-wave Newtonian noise
In this work, we evaluate the Newtonian noise (NN) arising from density perturbations induced by Rayleigh waves propagating on the ground surface, following the approximate formulation developed by Harms PhysRevD.88.122003 . Here we summarize the derivation from the vertical ground-displacement spectrum to the angular noise of a torsion-bar detector in a compact and self-contained manner.
Let $S_{x}^{\mathrm{ground}}(\Omega)\,[\mathrm{m}/\sqrt{\mathrm{Hz}}]$ denote the ASD of the vertical ground displacement associated with Rayleigh waves at angular frequency $\Omega$. The density fluctuations produced by near-surface Rayleigh waves generate horizontal gravitational perturbations acting on a test mass located at depth $d$ below the surface. The resulting displacement NN of a single test mass (in magnitude) can be written as
$S_{\rm R}(f)\simeq\frac{2\pi\,G\,\rho_{0}\,\gamma_{\rm R}}{\Omega^{2}}\,\langle|\cos\theta|\rangle_{\rm rms}\,S_{x}^{\mathrm{ground}}(\Omega)\,\exp\!\left(-\frac{\Omega d}{c_{\rm R}}\right),$ (36)
where $G$ is the gravitational constant, $\rho_{0}$ is the mean ground density, $c_{\rm R}$ is the Rayleigh-wave phase velocity, and $\gamma_{\rm R}$ is a dimensionless coupling factor that relates the ground displacement to the resulting gravitational perturbation. The angle $\theta$ is defined as the angle between the Rayleigh-wave propagation direction and the detector sensitivity axis. Assuming an isotropic distribution of propagation directions, we adopt
$\langle|\cos\theta|\rangle_{\rm rms}=\frac{1}{\sqrt{2}}.$ (37)
The exponential factor in Eq. (36) accounts for the attenuation of the Rayleigh-wave gravitational field with depth.
In the frequency range where the short-arm condition
$\frac{\Omega d}{c_{\rm R}}\ll 1$ (38)
is satisfied, the gravity-gradient (differential) displacement ASD between the two test masses can be approximated as
$S_{\rm GG,R}(\Omega)\simeq\mathcal{O}_{\rm R}\left(\frac{\Omega d}{c_{\rm R}}\right)S_{\rm R}(\Omega),$ (39)
where $\mathcal{O}_{\rm R}$ is a dimensionless orientation factor arising from angular averaging. For an isotropic distribution of Rayleigh-wave directions, we use
$\mathcal{O}_{\rm R}=\sqrt{\langle\cos^{4}\theta\rangle}=\sqrt{\frac{3}{8}}.$ (40)
Under the free-mass approximation, the corresponding differential acceleration ASD is given by
$\Delta a_{\rm R}(\Omega)=\Omega^{2}S_{\rm GG,R}(\Omega).$ (41)
For a torsion bar of length $L_{\mathrm{bar}}$ with point-like end masses $M$, the torque induced by this differential acceleration is
$\tau_{\rm R}(\Omega)=\left(\frac{L_{\mathrm{bar}}}{2}\right)M\Delta a_{\rm R}(\Omega).$ (42)
The mechanical susceptibility of the torsional mode is defined in Eq. 5. The angular ASD of the torsion bar induced by Rayleigh-wave Newtonian noise is then given by
$S_{\rm NN}(\Omega)$ $=$ $\eta_{g}I|\chi(\Omega)|\,\tau_{\rm R}(\Omega),$ (43)
$=$ $\frac{\eta_{g}}{\Omega^{2}}\tau_{\rm R}(\Omega).$ (44)
Table 7: Parameters used for the noise calcuation
Parameter Symbol Value
Gravitational constant $G$ $6.67430\times 10^{-11}\,\mathrm{m^{3}\,kg^{-1}\,s^{-2}}$
Boltzmann constant $k_{\mathrm{B}}$ $1.380649\times 10^{-23}\ \mathrm{J\,K^{-1}}$
Mean ground density $\rho_{0}$ $2000\,\mathrm{kg/m^{3}}$
Rayleigh coupling factor $\gamma_{\rm R}$ $0.83$
Rayleigh phase velocity $c_{\rm R}$ $3500\,\mathrm{m/s}$
Depth of observatory $d$ $200\,\mathrm{m}$
Angular average $\langle|\cos\theta|\rangle_{\rm rms}$ $1/\sqrt{2}$
Orientation factor $\mathcal{O}_{\rm R}$ $\sqrt{3/8}$
Young’s modulus $Y_{s}$ 425GPa
Poisson ratio $\sigma_{s}$ 0.1
Resonant frequency of ETM stage $\omega_{ETM}$ $2\pi\times 2.0\times 10^{-3}$
Internal mode frequency $\omega_{tb}$ $2\pi\times 667$
Pre-isolation stage gain $G_{pre}$ 1000
Temperature of ETM $T$ 8 K
Wire mechanical loss angle $\phi_{0}$ $1.0\times 10^{-7}$
Coating mechanical loss angle $\phi_{c}^{eff}$ $1.0\times 10^{-6}$
Substrate mechanical loss angle $\phi_{s}^{eff}$ $1.0\times 10^{-7}$